# A coisotropic embedding theorem for pre-multisymplectic manifolds

### v1

@@
A coisotropic embedding theorem for pre-multisymplectic manifolds
L. Schiavone1,2 [Uncaptioned image]
1Dipartimento di Matematica e Applicazioni Renato Caccioppoli, Università degli Studi di Napoli Federico II
2
e-mail: luca.schiavone@unina.it
Acknowledgements
I am indebted to prof. A. Ibort for making me aware of the existence of the coisotropic embedding theorem and its potential applications as well as for the huge amount of inspiring discussions of the last years. I am also indebted to prof. G. Marmo for encouraging me to think of alternative ways to depict the geometry behind Gotay’s coisotropic embedding theorem. I also want to thank prof. A. Zampini, F.M. Ciaglia and F. Di Cosmo for many useful discussions. I acknowledge financial support from Next Generation EU through the project 2022XZSAFN – PRIN2022 CUP: E53D23005970006. I am a member of the GNSGA (Indam).
1 Introduction
@@
Multisymplectic geometry has emerged as a fundamental framework for the geometric description of Field Theories, extending the role that symplectic geometry plays in Classical Mechanics, as testified by the huge, even if not exhaustive, list of contributions [GPR69, GPR71, Kij73, GS73, KS76, TK79, AA80, Got91a, Got91b, BSF88, GMS97, CCI91, FF03, EEMLRR96, EEMLRR00, RR09, Kru15, CDI+20b, CDI+20a, CDI+24, CDI+22b, CDI+22c] (see also references therein). In this setting, the tangent and cotangent bundles of a configuration space are replaced by the first-order jet bundle of a fibration and its dual, providing a natural arena for analyzing the equations of motion and symmetries of Field Theories. These jet bundles are examples of pre-multisymplectic manifolds, which generalize the notion of pre-symplectic geometry to higher-degree forms. In this framework, to any Lagrangian (or Hamiltonian), a pre-multisymplectic structure is associated, that allows for an intrinsic formulation of a variational principle providing the equations of motion of the theory.
@@
In previous contributions [CDI+20b, CDI+20a, CDI+24, CDI+22b, CDI+22c], we showed how, at least locally close to some Cauchy hypersurface, such equations of motion can be formulated in terms of (infinite-dimensional) pre-symplectic Hamiltonian systems. In particular, we established that the space of solutions of the equations of motion inherits a pre-symplectic structure. The kernel of this structure descends from the kernel of the pre-multisymplectic form mentioned above and encodes the gauge symmetries of the theory.
@@
To define a Poisson structure on the space of solutions, we employed a regularization technique based on M. Gotay’s coisotropic embedding theorem [Got82] (see also [GS90] for an equivariant version of the theorem and [OP05] for a more modern approach). Gotay’s theorem provides a powerful tool in symplectic geometry, ensuring that any pre-symplectic manifold can be embedded as a coisotropic submanifold into a symplectic manifold, referred to as its symplectic thickening. In [CDI+24, CDI+22b, CDI+22c], we used Gotay’s theorem to coisotropically embed the space of solutions of the equations of motion of Classical Field Theories into a symplectic manifold where a Poisson structure is canonically defined. We then studied whether and how this Poisson structure projects back to a Poisson structure on the space of solutions.
@@
Having in mind the previous statement that Field Theories can be locally formulated in terms of (infinite-dimensional) pre-symplectic Hamiltonian systems, the coisotropic embedding theorem has been used by the author also to provide a one-to-one correspondence between symmetries and constants of the motion [CDI+22a] and to provide a solution for the inverse problem of the calculus of variations for a class of implicit differential equations [Sch24b, Sch24a].
@@
The aim of this manuscript is to prove a coisotropic embedding theorem for pre-multisymplectic manifolds. Specifically, we will show that any pre-multisymplectic manifold can be embedded as a coisotropic submanifold into a larger multisymplectic manifold. This generalization provides a finite-dimensional framework for coisotropic regularization and opens new avenues for the study of multisymplectic geometry and its applications to Field Theories.
It is worth pointing out that recent work has focused on the interplay between coisotropic submanifolds and multisymplectic geometry, particularly in the opposite direction taken in this manuscript, namely that of multisymplectic reduction. In particular, in [de 24] the authors analyze coisotropic submanifolds of pre-multisymplectic manifolds and study conditions under which they inherit multisymplectic structures.
The structure of this paper is as follows. In Section 2, we recall preliminary notions about pre-multisymplectic manifolds that will be used throughout the manuscript. Section 3 is devoted to the statement and proof of the main theorem.
@@ 2 Preliminaries
A $k$-multisymplectic manifold is a smooth differential manifold $\mathcal{M}$ equipped with a closed and non-degenerate differential $k$-form $\omega$.
@@
A $k$-pre-multisymplectic manifold is a smooth differential manifold $\mathcal{M}$ equipped with a closed differential $k$-form $\omega$.
Definition 2.3 ($\ell$-multisymplectic orthogonal).
@@
[…]
@@
$\mathbf{T}_{n}\mathcal{N}^{\perp,\,\ell}\,=\,\mathbf{T}_{n}\mathcal{M}\,,\;\;\;\text{if }\,\,\ell>k\,.$ (3)
Definition 2.4 ($\ell$-Coisotropic submanifold).
@@
$\left\{\,q^{j}\,\right\}_{j=1,...,d}$ (4)
a system of local coordinates on it and by
@@
an adapted system of local coordinates on $\Lambda^{k}(\mathcal{M})$.
Consider a $k$-pre-multisymplectic manifold $(\mathcal{M},\,\omega)$. The kernel of $\omega$ is, at each $m\in\mathcal{M}$, a subspace of $\mathbf{T}_{m}\mathcal{M}$, denoted by $K_{m}$. We will always assume $\omega$ to have constant rank, so that $K_{m}$ and $K_{n}$ are isomorphic for any pair $m,\,n\in\mathcal{M}$. Moreover, we will also always assume that $K_{m}$ provides a completely integrable distribution on $\mathcal{M}$ [^1] Note that the complete integrability of the distribution generated by $K_{m}$ is ensured by the closure of $\omega$ only when $k=2$ by Frobenius theorem [AMR88]. so that there exists a foliation of $\mathcal{M}$ such that the tangent space to each leaf at each point $m\in\mathcal{M}$ coincides with $K_{m}$ and there exists a unique leaf passing through any point of $\mathcal{M}$. The space of leaves of the foliation is the space of equivalence classes of points on $\mathcal{M}$ belonging to the same leaf and will be denoted by
$\mathcal{M}/\sim_{K}\,.$ (6)
Sometimes, we will consider on $\mathcal{M}$ a system of local coordinates adapted to such a foliation
@@
where $l$ is the rank of $\omega$ and $r$ is the dimension of its kernel. The coordinates $x^{a}$ are a system of coordinates on the space of leaves of the foliation, namely, locally each leaf is a level set of the type
@@
whereas the coordinates $f^{A}$ (for fixed values of the $x^{a}$’s) individuates a point on each leaf. In this system of coordinates, $K_{m}$ reads
$K_{m}\,=\,\langle\left\{\,\frac{\partial}{\partial f^{A}}\,\right\}_{A=1,...,r}\rangle\,.$ (9)
@@
A connection on $\mathcal{M}$ is an idempotent smooth (1, 1)11(1,\,1)( 1 , 1 )-tensor field on $\mathcal{M}$ whose image, at each $m\in\mathcal{M}$, is $K_{m}$.
Locally a connection can be written as
@@
where the functions $P^{A}_{a}$ are the so-called connection coefficients (see, for instance [GMS10, App. $14$] or [GMS09, Sec. $1.3$]). By means of $P$, the tangent space to $\mathcal{M}$ at each point splits into the direct sum of tangent vectors in the image of $P$ and tangent vectors in the kernel of $P$. The image of $P$ reads
@@
$W_{m}\,=\,\langle\left\{\,W_{a}\,:=\,\frac{\partial}{\partial x^{a}}+P^{A}_{a}\frac{\partial}{\partial f^{A}}\,\right\}_{a=1,...,l;A=1,...,r}\rangle\,,$ (13)
and is usually referred to as the space of horizontal vectors. It is easy to show that under a change of local chart, the point-wise splitting into vertical and horizontal tangent vectors is preserved and, thus, the tangent bundle to $\mathcal{M}$ split as the direct sum
$\mathbf{T}\mathcal{M}\,=\,\mathbf{V}^{P}\oplus\mathbf{H}^{P}\,,$ (14)
where $\mathbf{V}^{P}$ (resp. $\mathbf{H}^{P}$) is the vector subbundle of $\mathbf{T}\mathcal{M}$ whose sections are, at each point, the vertical (resp. horizontal) vectors defined by $P$.
[…]
@@
the splitting of $\mathbf{T}\mathcal{M}$ induced by the connection $R$.
@@
$\alpha^{\parallel}_{P}\,:=\,\alpha(P(\,\cdot\,),...,P(\,\cdot\,))\,,$ (17)
@@
$\langle\left\{\,{\rm d}x^{a},\,{\rm d}f^{A}-P^{A}_{a}{\rm d}x^{a}\,=:\,\theta^{A}\,\right\}_{a=1,...,l;A=1,...,r}\rangle$ (19)
dual to the basis of vertical and horizontal vectors associated to $P$, one has
$\alpha\,=\,a_{j_{1}...j_{k}}\rho^{j_{1}}\wedge...\wedge\rho^{j_{k}}\,,$ (20)
where
$\begin{split}\rho^{j_{i}}\,&=\,{\rm d}x^{a}\,,\;\;\;\text{if}\;\;j_{i}=a\,,\\ \rho^{j_{i}}\,&=\,\theta^{A}\,,\;\;\;\text{if}\;\;j_{i}\,=\,A\,.\end{split}$ (21)
@@
$\begin{split}\alpha^{\parallel}_{P}\,&=\,\alpha_{A_{1}...A_{k}}\theta^{A_{1}}\wedge...\wedge\theta^{A_{k}}\,,\\ \alpha^{\perp}_{P}\,&=\,\sum_{I^{a}_{i}}\alpha_{I^{a}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}x^{a}}_{i}\wedge...\wedge\rho^{j_{k}}\,,\end{split}$ (22)
where $I^{a}_{i}$ is a multi-index $j_{1}...j_{k}$ where the $i$-th component is $a$ and the sum is taken over all the multi-indices of this type. We will denote by $\mathscr{J}$ the set of these multi-indices. As for vertical and horizontal tangent vectors, it is easy to show that this splitting of $k$-forms is stable under a change of local chart, and, thus, it defines two subbundles of $\Lambda^{k}(\mathcal{M})$. We will denote by
@@
$\alpha^{\parallel}_{R}\,:=\,\alpha(R(\,\cdot\,),...,R(\,\cdot\,))\,,$ (27)
@@
$\begin{split}\alpha^{\parallel}_{R}\,&=\,\alpha_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}\,,\\ \alpha^{\perp}_{R}\,&=\,\sum_{I^{A}_{i}}\alpha_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k}}\,,\end{split}$ (29)
where $I^{A}_{i}$ is a multi-index $j_{1}...j_{k}$ where the $i$-th component is $A$ and the sum is taken over all the multi-indices of this type. We will denote by $\mathscr{I}$ the set of these multi-indices.
@@
$\left\{\,q^{j},\,p_{I^{A}_{i}}\,\right\}_{j=1,...,d;I^{A}_{i}\in\mathscr{I}}\,,$ (32)
a system of local coordinates on ${\Lambda^{k}}^{\perp}_{P}(\mathcal{M})$.
3 The theorem
Here, we provide the statement and the proof of the multisymplectic coisotropic embedding theorem.
3.1 Statement

### v2

@@ Abstract
Motivated by applications in Geometric Field Theory, we prove a coisotropic embedding theorem à là Gotay [Got82] for pre-multisymplectic manifolds.
Introduction
@@
Multisymplectic geometry has emerged as a fundamental framework for the geometric description of Field Theories, extending the role that symplectic geometry plays in Classical Mechanics (see [TK79, BSF88, CCI91, RR09, Kru15] and references therein). In this setting, the tangent and cotangent bundles of a configuration space are replaced by the first-order jet bundle of a fibration and its dual, providing a natural arena for analyzing the equations of motion and symmetries of Field Theories. These jet bundles are examples of pre-multisymplectic manifolds, which generalize the notion of pre-symplectic manifold involving higher-degree forms. In this framework, to any Lagrangian (or Hamiltonian) function, a pre-multisymplectic structure is associated, which allows for an intrinsic formulation of a variational principle providing the equations of motion of the theory.
@@
In previous contributions [CDCI+20b, CDCI+20a, CDI+24, CDI+22b, CDI+23], we showed that, at least locally near a Cauchy hypersurface, the equations of motion can be formulated as (infinite-dimensional) pre-symplectic Hamiltonian systems. In this setting, the space of solutions inherits a pre-symplectic structure whose kernel reflects the kernel of the underlying pre-multisymplectic form, encoding the gauge symmetries of the theory.
@@
To define a Poisson structure on the space of solutions, we employed a regularization technique based on M. Gotay’s coisotropic embedding theorem [Got82] (see also [GS90] for an equivariant version of the theorem and [OP05] for a more modern approach). Gotay’s theorem provides a powerful tool in symplectic geometry, ensuring that any pre-symplectic manifold can be embedded as a coisotropic submanifold into a symplectic manifold, referred to as its symplectic thickening. In [CDI+24, CDI+22b, CDI+23], we used Gotay’s theorem to coisotropically embed the space of solutions of the equations of motion of Classical Field Theories into a symplectic manifold where a Poisson structure is canonically defined. We then studied whether and how this Poisson structure projects back to a Poisson structure on the space of solutions.
@@
Building on the observation that Field Theories can locally be expressed as (infinite-dimensional) pre-symplectic Hamiltonian systems, the author has also used the coisotropic embedding theorem to establish a one-to-one correspondence between symmetries and conserved quantities [CDI+22a], and to address the inverse problem of the calculus of variations for a class of implicit differential equations [Sch24b, Sch24a].
@@
The aim of this manuscript is to prove a coisotropic embedding theorem for pre-multisymplectic manifolds. Specifically, we will show that any pre-multisymplectic manifold can be embedded as a coisotropic submanifold into a larger multisymplectic manifold.
It is worth pointing out that recent work has focused on the interplay between coisotropic submanifolds and multisymplectic geometry, particularly in the opposite direction taken in this manuscript, namely that of multisymplectic reduction. In particular, in [dLIL24] the authors analyze coisotropic submanifolds of pre-multisymplectic manifolds and study conditions under which they inherit multisymplectic structures.
The structure of this paper is as follows. In Section˜1, we recall preliminary notions about pre-multisymplectic manifolds that will be used throughout the manuscript. Section˜2 is devoted to the statement and proof of the main theorem, and Section˜3 provides an example to illustrate the main construction of the theorem.
@@ 1 Preliminaries
$\left\{\,q^{j}\,\right\}_{j=1,...,d}$ (1)
a system of local coordinates on it.
Recall the following definitions.
@@
A $k$-plectic manifold (or multisymplectic manifold) is a smooth differential manifold $\mathcal{M}$ equipped with a closed and non-degenerate differential $k$-form $\omega$.
@@
A pre-$k$-plectic manifold (or pre-multi-symplectic manifold) is a smooth differential manifold $\mathcal{M}$ equipped with a closed differential $k$-form $\omega$.
Definition 1.3 ($\ell$-multisymplectic orthogonal [CIDL99]).
@@
$\mathbf{T}_{n}\mathcal{N}^{\perp,\,\ell}\,=\,\left\{\,V\in\mathbf{T}_{n}\mathcal{M}\;\;:\;\;i_{V\wedge W_{1}\wedge...\wedge W_{\ell}}\omega\,=\,0\,,\;\;\;\forall\,\,W_{1},\,...,\,W_{\ell}\in\mathbf{T}_{n}\mathcal{N}\,\right\}\,$, (2)
where, with a slight abuse of notation, the tangent vectors $W_{1},\,...,\,W_{\ell}$ in the contraction above denote the tangent vectors $\mathfrak{i}_{\star}W_{1},\,...,\,\mathfrak{i}_{\star}W_{\ell}$, $\mathfrak{i}$ denoting the canonical embedding of $\mathcal{N}$ into $\mathcal{M}$.
@@
$\mathbf{T}_{n}\mathcal{N}^{\perp,\,\ell}\,=\,\mathbf{T}_{n}\mathcal{M}\,,\;\;\;\text{if }\,\,\ell>\operatorname{dim}\mathbf{T}_{n}\mathcal{N}\,.$ (4)
Definition 1.4 ($\ell$-Coisotropic submanifold [CIDL99]).
@@
We will usually denote by
@@
a system of local coordinates on $\Lambda^{k}(\mathcal{M})$ adapted to the system of coordinates on $\mathcal{M}$ mentioned above, where the fibered coordinates $p_{j_{1}...j_{k}}$ have to be understood as the $k$-forms
$p_{j_{1}...j_{k}}{\rm d}q^{j_{1}}\wedge...\wedge{\rm d}q^{j_{k}}\,.$ (6)
Consider a pre-$k$-plectic manifold $(\mathcal{M},\,\omega)$. The kernel of $\omega$ is, at each $m\in\mathcal{M}$, a subspace of $\mathbf{T}_{m}\mathcal{M}$, denoted by $K_{m}$. We will always assume $\omega$ to have constant rank, so that $K_{m}$ and $K_{n}$ are isomorphic for any pair $m,\,n\in\mathcal{M}$. Given that $\omega$ is closed and of constant rank, by means of Frobenius theorem one shows that $K_{m}$ provides a completely integrable distribution on $\mathcal{M}$, so that there exists a foliation of $\mathcal{M}$ such that the tangent space to each leaf at each point $m\in\mathcal{M}$ coincides with $K_{m}$ and there exists a unique leaf passing through any point of $\mathcal{M}$. We refer to [CC00] for the basic notions about foliations we will use along the manuscript. Sometimes, we will consider on $\mathcal{M}$ a system of local coordinates adapted to such a foliation
@@
where $l$ is the rank of $\omega$ and $r$ is the dimension of its kernel, so that $l+r=d$. Coordinates $x^{a}$ are a system of coordinates on the space of leaves of the foliation, namely, locally each leaf is a level set of the type
@@
whereas coordinates $f^{A}$ (for fixed values of the $x^{a}$’s) individuate a point on each leaf. In this system of coordinates, $K_{m}$ reads
where the symbol $\langle\,\cdot\,\rangle$ denotes the linear span of the vectors appearing inside the angle brackets.
@@
A connection on $\mathcal{M}$ is an idempotent [^1] By idempotent we mean that $P^{2}=P$. smooth (1, 1)(1,\,1)-tensor field on $\mathcal{M}$ whose image, at each $m\in\mathcal{M}$, is $K_{m}$.
Locally, a connection can be written as
@@
where the functions $P^{A}_{a}$ are the so-called connection coefficients. By means of $P$, the tangent space to $\mathcal{M}$ at each point splits into the direct sum of tangent vectors in the image of $P$ and tangent vectors in the kernel of $P$. Being a (1, 1)(1,\,1)-tensor field on $\mathcal{M}$, $P$ defines a map from vector fields on $\mathcal{M}$ to vector fields on $\mathcal{M}$ via contraction. The image of such a map reads
$K_{m}\,=\,\langle\left\{\,V_{A}\,:=\,\frac{\partial}{\partial f^{A}}\biggr|_{m}\,\right\}_{A=1,...,r}\rangle\,$, (12)
@@
$H_{m}\,=\,\langle\left\{\,H_{a}\,:=\,\frac{\partial}{\partial x^{a}}\biggr|_{m}+P^{A}_{a}(m)\frac{\partial}{\partial f^{A}}\biggr|_{m}\,\right\}_{a=1,...,l}\rangle\,$, (13)
and is usually referred to as the space of horizontal vectors. The pointwise splitting $\mathbf{T}_{m}\mathcal{M}=K_{m}\oplus H_{m}$ extends to a global splitting of the tangent bundle $\mathbf{T}\mathcal{M}=\mathbf{V}^{P}\oplus\mathbf{H}^{P}$ because both $\mathbf{V}^{P}$ and $\mathbf{H}^{P}$ are smooth vector sub-bundles. The vertical bundle $\mathbf{V}^{P}$ corresponds to the kernel of the pre-multisymplectic form $\omega$. Since $\omega$ is a smooth form assumed to have constant rank, its kernel is a well-defined smooth sub-bundle. The horizontal bundle $\mathbf{H}^{P}$ is defined as the kernel of the connection $P$. By construction, $P$ is a smooth (1, 1)(1,\,1)-tensor field. As a projector onto a constant-rank bundle, $P$ also has constant rank, and thus its kernel $\mathbf{H}^{P}$ is also a smooth vector sub-bundle. The direct sum of two smooth sub-bundles that span the tangent bundle at every point constitutes a global vector bundle splitting. Sections of $\mathbf{V}^{P}$ (resp. $\mathbf{H}^{P}$) are called vertical (resp. horizontal) vector fields.
Given any connection $P$, the related (1, 1)(1,\,1)-tensor field $R\,=\,\mathbf{1}-P$, where $\mathbf{1}$ denotes the identity operator, provides a complementary splitting to that induced by $P$, where the horizontal space coincides with the vertical space of $P$ and vice versa. We will denote by
@@
the splitting of $\mathbf{T}\mathcal{M}$ induced by $R$.
@@
$\alpha^{\parallel}_{P}\,:=\,\alpha(\underbrace{P(\,\cdot\,),...,P(\,\cdot\,)}_{\text{k times}})\,$, (16)
@@
$\left\{\,{\rm d}x^{a},\,{\rm d}f^{A}-P^{A}_{a}{\rm d}x^{a}\,=:\,\theta^{A}\,\right\}_{a=1,...,l;A=1,...,r}\,$, (18)
dual to the basis of vertical and horizontal vector fields associated with $P$, namely such that
[…]
one has
$\begin{split}\alpha\,=\,&\alpha_{A_{1}...A_{k}}\theta^{A_{1}}\wedge...\wedge\theta^{A_{k}}+\\ &\alpha_{a_{1}A_{2}...A_{k}}{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge...\wedge\theta^{A_{k}}+\\ &\;\;\vdots\\ &\alpha_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}\,.\end{split}$ (20)
@@
$\alpha^{\parallel}_{P}\,=\,\alpha_{A_{1}...A_{k}}\theta^{A_{1}}\wedge...\wedge\theta^{A_{k}}\,$, (21)
and
$\alpha^{\perp}_{P}\,=\,\alpha_{a_{1}A_{2}...A_{k}}{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge...\wedge\theta^{A_{k}}+...+\alpha_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}\,.$ (22)
Since this decomposition is defined pointwise using the smooth tensor field $P$, it is itself smooth. Building upon the established global splitting of the tangent bundle, this construction naturally defines a splitting of the bundle of $k$-forms into two smooth vector sub-bundles. We will denote by
@@
$\alpha^{\parallel}_{R}\,:=\,\alpha(\underbrace{R(\,\cdot\,),...,R(\,\cdot\,)}_{\text{k times}})\,$, (27)
@@
$\alpha^{\parallel}_{R}\,=\,\alpha_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}\,$, (29)
while
$\begin{split}\alpha^{\perp}_{R}\,=\,&\alpha_{A_{1}...A_{k}}\theta^{A_{1}}\wedge...\wedge\theta^{A_{k}}+\\ &\alpha_{a_{1}A_{2}...A_{k}}{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge...\wedge\theta^{A_{k}}+\\ &\;\;\vdots\\ &\alpha_{a_{1}...a_{k-1}A_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k-1}}\wedge\theta^{A_{k}}\,.\end{split}$ (30)
@@
$\left\{\,q^{j},\,p_{A_{1}...A_{k}},\,p_{a_{1}A_{2}...A_{k}},\,...,\,p_{a_{1}...a_{k-1}A_{k}}\,\right\}_{j=1,...,d;a_{k}=1,...,r;A_{k}=1,...,l}\,$, (33)
a system of local coordinates on ${\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$.
2 Coisotropic embeddings of pre-multisymplectic manifolds
We are now ready to prove a coisotropic embedding theorem à là Gotay for pre-multisymplectic manifolds.